a random sample of 90 observations produced a mean of xbar = 29.3 and s = 3.8. Find an approximate 93% confidence interval for the mean.
Answer #1

Solution :

Given that,

Point estimate = sample mean = = 29.3

sample standard deviation = s = 3.8

sample size = n = 90

Degrees of freedom = df = n - 1 = 89

t /2,df = 1.834

Margin of error = E = t/2,df * (s /n)

= 1.834 * (3.8 / 90)

Margin of error = E = 0.7

The 93% confidence interval estimate of the population mean is,

- E < < + E

29.3 - 0.7 < < 29.3 + 0.7

28.6 < < 30.0

(28.6 , 30.0)